OpenStudy (anonymous):

How we find equation of ellipse with given data,when given, 1.foci and minor axis of length2.foci and major axis of length and foci and vertices?any link for this thanks !!!!

OpenStudy (anonymous):

i have many questions on it if u give me link from where i get the way to solve those questions

OpenStudy (anonymous):

am its too much easy ..... wats the general equation for ellipse ... write its down here

OpenStudy (anonymous):

no i hve to make equation from its elements

OpenStudy (anonymous):

i.e ellipse

OpenStudy (sirm3d):

 equation of ellipse is \[\large\frac{ x^2 }{ a^2 }+\frac{y^2}{b^2}=1\] minor axis: b, major axis: a Foci: \[(\large \pm c,0)\] where \[\large c^2 = a^2 - b^2\]

OpenStudy (sirm3d):

the relation between (lengths of) minor axis, major axis and (coordinates of) foci is determined by \[\huge c^2 = a^{2} - b^2\]

OpenStudy (anonymous):

sirm3d m confused can u solve this ,find an equation of ellipse with foci + and _3,0 and minor axis of length 10?

OpenStudy (sirm3d):

c=3, b=10. solve a.

OpenStudy (anonymous):

a sq=34

OpenStudy (sirm3d):

use these values of a and b and you have the equation of the ellipse

OpenStudy (anonymous):

yeah i got it thanks but how we get c which is equal to 3?

OpenStudy (sirm3d):

the value of a^2 you gave me is not right, though.

OpenStudy (sirm3d):

foci at (+/- c,0) and you gave me the foci at (+/- 3,0), so c = 3.

OpenStudy (anonymous):

what is c?

OpenStudy (anonymous):

u know m very dull in mathematis

OpenStudy (sirm3d):

\[\huge (\pm c,0) = (\pm 3,0)\]. now you tell me the value of c.

OpenStudy (anonymous):

3

OpenStudy (sirm3d):

right.

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

c2=a2-b2 now solve it urslf ... c focus, a major axis, b minur axis. find the values and then put in its standard equation of ellips

OpenStudy (anonymous):

\now get it from pic

OpenStudy (anonymous):

khuram i have already solved equation ll be x sq\34 + ysq\25 m i right?

OpenStudy (anonymous):

c=3 and 2b=10 imlies that b=5 using csq=asq-bsq(sq=square) which give us asq=34 now required equation ll be xsq divided by asq + ysq divided by bsq

OpenStudy (anonymous):

daer actaully u didn`t mention the origion corrdinates.. so i can`t guess axactly and u didnt mention here taht which is major and which one is minor exis ... can u send me origional question

OpenStudy (anonymous):

just puting values of a sq and b sq we ll get an equation

OpenStudy (anonymous):

yeah wait..

OpenStudy (anonymous):

yes its ryt .. ryt now i got ur question a little more clear now

OpenStudy (anonymous):

foci\[\pm3,0 and minor axis of length 10\]

OpenStudy (anonymous):

i have solved it just check it whether it's right or not ok

OpenStudy (anonymous):

yes i got it its origion iz (0,0) yah its ryt mam.

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

yeah it is reallly easy

OpenStudy (anonymous):

no need to pay thanks u did it uself i apprecaite ur attemp ... welldone

OpenStudy (anonymous):

i was wondering what is c,it was creating problems for me but from ur pdf i came to know it si foci heh:)

OpenStudy (anonymous):

ok welll .. u got ur question tahst great